OpenStudy (anonymous):

Simplify the following expression. √-121√-24

OpenStudy (anonymous):

\[\sqrt{-121*-24}=\sqrt{121*24}=\sqrt{11*11*4*6}=11*2*\sqrt{6}=22\sqrt{6}\]